# OViP: Online Vision-Language Preference Learning

### v1

@@ 3.4 Ablation Study
We evaluated various combinations of loss functions for online preference learning in hallucination mitigation to derive the final formulation in Equation 6. Our ablation study examines the effectiveness of different training objectives, including text-side ($\mathcal{L}_{\mathrm{Text}}$), image-side and auxiliary losses. Specifically for image-side losses, we examine our image loss $\mathcal{L}_{Image}$ alongside two variants $\mathcal{L}_{Image}^{base}$ and $\mathcal{L}_{Image\mathrm{-}Sym}$. For auxiliary loss, we compare the anchor loss proposed by [19] and the bidirectional anchor loss, which enforce the probability of positive response to increase and the negative one to decrease. Detailed formulations are provided in Appendix B.1.
Table 2: Results of different loss functions on hallucination-related benchmarks, with key metrics shown. OViP loss is $\mathcal{L}_{\mathrm{Text}}+\mathcal{L}_{\mathrm{Image}}$. “--” means the removal of the corresponding loss from OViP.
Loss Functions Training from scratch Iterative training
AMBgen MMHal AMBdis HRI ObjectHal MMHal AMBdis HRI
OViP 65.38 2.35 86.6 4.32 74.14 2.70 85.7 7.94
      $-\ \mathcal{L}_{\mathrm{Text}}$ 66.61 2.12 85.4 4.23 74.01 2.60 86.1 7.71
      $-\ \mathcal{L}_{\mathrm{Image}}$ 63.58 2.06 85.9 -2.29 72.63 2.40 86.6 4.56
$\mathcal{L}_{\mathrm{Text}}+\mathcal{L}_{\mathrm{Image}}^{\ base}$ 65.10 2.17 86.2 4.08 73.45 2.58 86.3 7.50
$\mathcal{L}_{\mathrm{Image-Sym}}$ 65.54 2.09 85.6 -0.32 73.86 2.41 86.3 6.57
@@
Figure 3: Effect of auxiliary loss on text-side and image-side losses. HRI measures hallucination-reduction, while General Score reflects general performance change. Marker color indicates the primary loss function used, and marker shape denotes the usage of auxiliary loss.
Figure 4: Results of offline and online training strategy with DPO and OViP. Cover measures the informativeness of model response from AMBER generative benchmark.
We further analyze the impact of auxiliary loss in Figure 4. Contrary to the findings in [19], we find that incorporating anchor loss consistently reduces general capability and increases hallucinations across all other loss combinations. Moreover, while applying bi-directional anchor loss slightly improves general capabilities, it does not necessarily enhance hallucination mitigation. Among them, the OViP loss without any auxiliary loss achieves the best performance on both hallucination reduction and general capabilities. We therefore adopt OViP loss as our default training loss function.
@@
Figure 4 demonstrates that online training significantly outperforms its offline counterpart. For the offline setting, we construct training data by sampling and filtering model outputs prior to training. Whether using DPO or our proposed OViP loss, online training consistently outperforms offline training by a margin of no less than 4 points on the HRI metric within just one epoch. This improvement is not limited to the aggregated metric: across every individual benchmark and each corresponding metric, online training yields more stable and superior performance. Detailed results are provided in Appendix B.3.
Another notable observation is that online training also enhances the informativeness of model outputs. Even when trained solely with DPO, the Cover score remains above 50. In contrast, previous studies [16, 28, 21] using similar datasets typically exhibit a drop in this aspect.
@@ 4 Further Study
Table 3: Experiments about the training process.
(a): Token-wise average logprobability. Base DPOOff DPOOn IND -0.90 -0.97 -1.08 OOD -1.23 -1.17 -1.17
(b): Sampling with $temp.=1.2$, $num\_return\_sequences=16$ Steps 0 10 100 500 Num. 14.4 15.2 15.3 15.5
(c): Negative image construction R.Crop Offline Online HRI 9.76 9.23 10.88 General 0.34 -0.88 1.44
@@
Figure 5: Results using different training strategy. Offline represents training with existing dataset. Off Policy means training with sampled data and Iterative means the dataset of the epoch 2 is from model sampling.
Figure 6: Comparison of response quality distributions after training with different methods. The x-axis represents the semantic quality of model responses, with the leftmost region indicating severe hallucinations. The y-axis denotes the probability density over sampled responses. A rightward shift in the distribution corresponds to higher overall response quality and reduced hallucination.
Online Learners Gets More
Figure 6 illustrates how hallucination-related metrics evolve during iterative training under different optimization strategies. A key finding is that all non-online methods exhibit varying degrees of overfitting in multi-round training, as evidenced by a decline in the HRI. In contrast, only online strategies consistently yield stable improvements. Notably, training with external datasets leads to severe overfitting in the second round. Some prior works (such as [28, 21]) adopt at least two rounds of training on small, static datasets. While they report significant gains on certain metrics—such as AMBgen Chair and ObjectHal Chairr—a more holistic view reveals substantial overfitting. In Table 1, HA-DPO, RLAIF-V, and OPA-DPO fall into this category.
Both Online DPO and Off-Policy DPO exhibit an initial drop in performance, meanwhile, GRPO and OViP show relatively slow improvement during the early stages of training. This phenomenon is attributed to the model’s initially skewed output distribution, which has a low perplexity over few responses it generates (Table 3(a)(b)). At this stage, the model tends to forget more than it learns, leading to a temporary decline in performance. As training progresses and the distribution gradually smooths out, the model begins to receive more diverse and informative supervision, allowing it to shift toward higher-score regions.
In contrast, Off-Policy DPO continues to rely on fixed data sampled from the early-stage distribution. As the model evolves, this static data becomes increasingly misaligned with the model’s current distribution, leading to ineffective learning due to the mismatch between the training samples and the model’s evolving behavior.
Learning Dynamics
To investigate output distribution shifts, we focus on hard queries in the test set and compare the distributions of sampled responses across models. Our key findings are: (1) Online learning enables both hallucination suppression and promotion of high-quality responses, effectively optimizing both ends of the distribution; (2) Image contrastive learning leads to overall quality improvement. Moreover, as shown in Table 3(c), the quality of negative image samples also matters. Hard negatives constructed via online strategies yield the best training results.
@@ 5.1 LVLM Hallucination
A wide range of prior work has explored synthetic data construction for mitigating hallucination in VLMs, which can be broadly categorized into image-related synthesis and text-only synthesis. On the image side, several approaches leverage entity extraction and masking to perform targeted image editing, generating visually similar but semantically distinct counterfactuals [17, 23]. In contrast, HallusionBench [42] adopts a manual approach, carefully crafting counterfactual images to probe specific failure modes. Other works take a generative perspective: SynthVLM [43] and SynthEmbedding [44] utilize off-the-shelf models to synthesize new images or directly generate image embeddings for hallucination-aware training.
Meanwhile, a larger body of research focuses on text-side data augmentation. VoCoT [45] introduces new prompting patterns and response types to generate hallucination-prone QA data at scale. Other works such as [22, 18, 33] introduce noise through perturbation, masking, or controlled corruption to simulate erroneous responses. More recent approaches [40, 15] aim to detect and correct hallucinated content at varying levels of granularity, from token-level edits to full-sequence rewrites.
@@
5.2 Allocating Computation on Training Sample Construction
Recent research has increasingly adopted the paradigm of allocating additional computation during training to get better training samples. Several studies utilize reinforcement learning with human or AI-generated feedback to guide VLM outputs. RLHF-V [15] leverages fine-grained human annotations to correct hallucinated content, while RLAIF-V [16] replaces human labels with feedback from ensembles of open-source models, significantly reducing annotation overhead. Similarly, OPA-DPO [28] employs an on-policy editing step prior to DPO, aligning training samples closely with model predictions to enhance data efficiency. Active learning methods further embody this approach, selectively querying uncertain or diverse samples to maximize training informativeness [46], while CLIP-based methods dynamically filtered self-generated samples for high-quality training pairs [47]. Other methods integrate auxiliary reward models or evaluators during training, providing continuous and adaptive feedback loops [32, 48]. Finally, recent approaches incorporate reasoning or editing mechanisms directly into training, using iterative self-feedback or generative data augmentation techniques to dynamically refine model outputs [31, 49]. These strategies improve model alignment and factuality by enriching the quality and relevance of supervision signals during training.
@@ 6 Conclusion
In this work, we propose the Online Vision-language Preference Learning (OViP) framework to efficiently address the pervasive hallucination problem in LVLMs. By integrating online preference learning with image-aware training, OViP enables real-time construction of high-quality contrastive data during training—without relying on a teacher LVLM. Furthermore, to better assess the trade-offs between hallucination reduction and overall performance, we refine and extend existing evaluation protocols, proposing a more comprehensive metric suite. Experimental results demonstrate that OViP significantly outperforms prior offline/online training approaches, achieving substantial hallucination reduction while preserving general vision-language capabilities. Our investigation into training-time dynamics also sheds light on the underlying mechanisms behind OViP’s effectiveness.
@@ A.1.1 Metric Design
[…]
@@ A.1.5 Further Study
A.2 Bad Cases
A.2.1 MMHal
@@
Figure 9: OPA-DPO fails to mention the man, a deficiency that is captured by the “Cover” metric but often overlooked in previous evaluations. “vehicles” is incorrectly identified as a hallucination word.
A.2.2 AMBER-generative & ObjectHal
@@ B.2.2 Ablation on Online Learning
B.2.3 Further Study
We save several checkpoints during training and evaluate them to gain a complete understanding of the training process.
We select 227 instances from the original OPA-DPO dataset—distinct from the training set—for analyzing changes in token-wise log-probability and output quality distributions.
For token-wise log-probability analysis, we treat the original negative samples from the dataset as out-of-distribution (OOD) responses and compute their log-probabilities. For in-distribution (IND) responses, we perform 16 response samplings per query using our model with a temperature of 0.2, and compute the average log-probability across these samples.
To examine output distribution shifts, we sample 16 responses per query with a higher temperature of 1.2, assign scores to each response, and analyze the distribution of these scores.
For comparison about different image generation strategy, we implement two other representative methods. 1. Random Cropping (R.Crop). We randomly crop 02̃0% area from the image to form the negative image. This method is from mDPO. 2. Offline construction. We use an LLM to generate the image description with some inaccordance with the positive image, then we use the diffusion model to generate the image according to the description. The prompt is at Table 8.
@@ Appendix E Limitations
Appendix F Broader Impacts
Appendix G Prompts for Judgment and Negative Image Generation
Prompt for Quality Judgment # Task
Your role is as a discerning assistant tasked with evaluating model responses for multimodal tasks (though you have no access with the image). Upon being presented with a question that requires the interpretation of both text and images, you will receive two distinct responses. The first is crafted by our sophisticated multimodal model, while the second represents an approximate ideal answer–it may be incomplete. Your objective is to meticulously and precisely assess the model-generated response (the former) based on the provided reference answer (the latter).
- Here’s how you should approach the assessment process:
11111. The quality of the response depends on its accuracy and the degree of adherence to the correct answer. Therefore, if the response is much more detailed than the reference answer, it should not be considered a very good response (although it may still be considered a good one).
@@ Appendix G Prompts for Judgment and Negative Image Generation
11113. Please do not provide additional reasoning, just give the score directly.
# Question
{question}
# Response
{response}
# Correct answer
{answer}
@@
Prompt for image generation-Part 1 # Task
Given an unknown image-related question, a correct answer, and an inaccurate response, carefully analyze the differences between the response and the answer. Then, provide a brief description of the image so that it aligns with the correct answer and differs from the incorrect responses. In other words, infer the content of the image.
# Example
[Example 1]
**<Question>**
What is on the sandwich?
**<Answer>**
The sandwich has tomatoes and lettuce on it.
**<Response>**
The sandwich has a slice of egg and tomato on it.
**<Output Description>**
A sandwich with only tomatoes and lettuce on it.
**Explanation**: The answer mentions lettuce and tomato, while the incorrect response mentions tomato and egg. So there is no egg on the sandwich.
Prompt for image generation-Part 2 [Example 2]
**<Question>**
Can you point out the details that make this image unique?
**<Answer>**
In the image, there is a plate with a slice of pizza topped with tomatoes, herbs, and cheese. The distinctive detail about the image is that the pizza is missing two slices, leaving just one slice remaining on the plate. This suggests that someone has already started enjoying the pizza. The slice appears to be well-cooked and freshly served, creating an appetizing and mouthwatering scene for the viewer.
**<Response>**
In the image, there is a slice of pizza on a plate with tomatoes and cheese. The pizza appears to be homemade and has been cut into two pieces. The tomatoes are sliced in half, revealing their juicy interior. The cheese on top of the pizza is melted, creating a delicious-looking dish. Additionally, there is a fork nearby, suggesting that someone might be planning to enjoy this pizza soon.
**<Output Description>**
A plate with a one-third remaining piece of pizza, topped with herbs, cheese, and tomatoes; someone has finished eating and left.
**Explanation**: The answer mentions that only one-third of the pizza remains and that someone has just finished eating and left, which is inconsistent with the response. Therefore, the image should include these two features.
[Example 3]
**<Question>**
Bird or cow?
**<Answer>**
Bird
**<Response>**
The bird in the image is a small, brown and white bird with a distinctive head shape and coloration. It is not a cow. The bird is perched on a branch, which is situated in front of a white building.
**<Output Description>**
A big, blue bird perched on a branch in front of a black building.
**Explanation**: Both the answer and the response mention the bird, but the response is more detailed. So the description should be contrastive to the features of the bird in the response.
# Requirements
- The description should be brief but precise.
- If both the answer and the response are long, focus on describing the one or two most significant differences.
- Do not provide any analysis or explanation; only describe the image.
- A common approach is to describe what is present in the image and what is missing.
**<Question>**
{question}
**<Answer>**
{answer}
**<Response>**
{response}
**<Output Description>**
Table 6: The prompt for image generation instruction.
Prompt for image distortion-Part 1 # Task
Given an unknown image-related question, a correct answer, and an inaccurate response, carefully analyze the differences between the response and the answer. Then, provide a brief description of the image so that it aligns with the correct answer and differs from the incorrect responses. In other words, infer the content of the image.
# Example [Example 1]
**<Question>**
What is on the sandwich?
**<Answer>**
The sandwich has tomatoes and lettuce on it.
**<Response>**
The sandwich has a slice of egg and tomato on it.
**<Output Description>**
A sandwich with only tomatoes and lettuce on it.
**Explanation**: The answer mentions lettuce and tomato, while the incorrect response mentions tomato and egg. So there is no egg on the sandwich.
[Example 2]
**<Question>**
Can you point out the details that make this image unique?
**<Answer>**
In the image, there is a plate with a slice of pizza topped with tomatoes, herbs, and cheese. The distinctive detail about the image is that the pizza is missing two slices, leaving just one slice remaining on the plate. This suggests that someone has already started enjoying the pizza. The slice appears to be well-cooked and freshly served, creating an appetizing and mouthwatering scene for the viewer.
**<Response>**
In the image, there is a slice of pizza on a plate with tomatoes and cheese. The pizza appears to be homemade and has been cut into two pieces. The tomatoes are sliced in half, revealing their juicy interior. The cheese on top of the pizza is melted, creating a delicious-looking dish. Additionally, there is a fork nearby, suggesting that someone might be planning to enjoy this pizza soon.
**<Output Description>**
A plate with a one-third remaining piece of pizza, topped with herbs, cheese, and tomatoes; someone has finished eating and left.
**Explanation**: The answer mentions that only one-third of the pizza remains and that someone has just finished eating and left, which is inconsistent with the response. Therefore, the image should include these two features.
Table 7: The prompt for distorted image generation instruction.
Prompt for image distortion-Part 2 [Example 3]
**<Question>**
Bird or cow?
**<Answer>**
Bird
**<Response>**
The bird in the image is a small, brown and white bird with a distinctive head shape and coloration. It is not a cow. The bird is perched on a branch, which is situated in front of a white building.
**<Output Description>**
A big, blue bird perched on a branch in front of a black building.
**Explanation**: Both the answer and the response mention the bird, but the response is more detailed. So the description should be contrastive to the features of the bird in the response.
# Requirements
- The description should be brief but precise.
- If both the answer and the response are long, focus on describing the one or two most significant differences.
- Do not provide any analysis or explanation; only describe the image.
- A common approach is to describe what is present in the image and what is missing.
**<Question>**
question
**<Answer>**
answer
**<Response>**
response
**<Output Description>**
Table 8: The prompt for distorted image generation instruction.

### v2

@@ 3.4 Ablation Study
We evaluated various combinations of loss functions for online preference learning in hallucination mitigation to derive the final formulation in Equation 6. Our ablation study examines the effectiveness of different training objectives, including text-side ($\mathcal{L}_{\mathrm{Text}}$), image-side and auxiliary losses. Specifically for image-side losses, we examine our image loss $\mathcal{L}_{Image}$ alongside two variants $\mathcal{L}_{Image}^{base}$ and $\mathcal{L}_{Image\mathrm{-}Sym}$. For auxiliary loss, we compare the anchor loss proposed by Wang et al. (2024a) and the bidirectional anchor loss, which enforce the probability of positive response to increase and the negative one to decrease. Detailed formulations are provided in Appendix B.1.
Table 2: Results of different loss functions. $\mathcal{L}_{\mathrm{OViP}}=\mathcal{L}_{\mathrm{Text}}+\mathcal{L}_{\mathrm{Image}}$.
Loss Functions HRI
From Scratch Iterative
$\mathcal{L}_{\mathrm{OViP}}$ -4.32 7.94
@@
[…]
$\mathcal{L}_{\mathrm{Text}}+\mathcal{L}_{\mathrm{Image}}^{\ base}$ -4.08 7.50
$\mathcal{L}_{\mathrm{Image-Sym}}$ -0.32 6.57
Table 3: Results of offline and online training strategy with DPO and OViP. Cover measures the informativeness of the model from AMBgen. Cover score of the original model is 50.
Method Cover HRI General AccDif
OViP online 51.1 -9.36 -0.88
offline 49.9 -4.32 -0.08
DPO online 50.0 -1.71 -2.57
offline 48.3 -2.29 -1.38
[…]
Figure 3: Effect of applying auxiliary loss to different loss functions.
Under the same iterative training regime, we further analyze the effect of auxiliary losses based on the DPO-initialized model and its sampled responses, as illustrated in Figure 3. Contrary to the findings in mDPO (Wang et al., 2024a), we find that incorporating anchor loss consistently reduces general capability and increases hallucinations across all loss combinations. Moreover, while applying bi-directional anchor loss slightly improves general capabilities, it does not necessarily enhance hallucination mitigation. Therefore, OViP loss without anchor loss is the most effective training objective for both reducing hallucination and maintaining general ability.
@@
Table 3 demonstrates that online training consistently outperforms its offline counterpart in HRI by at least 4 points within just one epoch (and continues to improve with further training, while offline training suffers from overfitting). Another notable observation is that online training also improves the informativeness of model outputs. Even when trained with DPO, the Cover score remains 50. In contrast, previous studies (Yu et al., 2024b; Yang et al., 2025; Fu et al., 2025) using the similar dataset typically exhibit a drop in this aspect. Additionally, the improvement for online training over offline training is almost across every individual benchmark and each corresponding metric, online training yields more stable and superior performance. Detailed results are provided in Appendix B.3.
@@ 4 Further Study
4.1 Training Efficiency
[…]
@@ 4.1 Training Efficiency
Figure 5: Results under different training strategies. Offline denotes training with the existing dataset; Off Policy refers to training with sampled data; and Iterative indicates that the dataset for the second epoch is generated by sampling from the offline-trained model after the first epoch.
[…]
4.2 Training Dynamics
Figure 5 illustrates how HRI evolves during training under different strategies, which allows us to investigate the dynamics of hallucination throughout training.
Need for Visual and Online Signals
For hallucination mitigation in LVLMs, adding visual supervision signals proves crucial: offline OViP surpasses GRPO and all DPO variants with one epoch. Building on this, online methods offer further advantages, which not only make each optimization step more effective in reducing hallucinations, but also exhibit better scalability, with overfitting arising significantly later compared to non-online approaches, whose performance starts to drop after training for one epoch. We conjecture that this superiority stems from the model-specific nature of hallucinations, which requires supervision to precisely target the current model’s errors.
Early Training Stagnation
Both Online DPO and Off-Policy DPO exhibit an initial drop in performance, while GRPO and OViP show relatively slow improvement during the early stages of training. We attribute this phenomenon to the model’s initially skewed output distribution. Early training primarily increases the diversity of model outputs, which does not immediately translate into performance gains but expands the search space for subsequent learning. A detailed discussion is provided in Appendix B.4.
Figure 6: Change in probability mass for the responses with corresponding score after training. “Low score” refers to scores less than 4. We smooth the discrete probability changes over 11 score bins (0-10) into a continuous curve. “Change > 0” represents the probability increases after training.
4.3 What Do We Actually Optimize during Training?
We investigate how training affects the distribution of response quality, as shown in Figure 6. The results indicate that performance gains come from both reducing hallucinated outputs and increasing high-quality ones. After training, responses with scores below 6 become less likely, while those above 6 occur more frequently.
Online training yields additional benefits over offline DPO. It further lowers the probability of severely hallucinated responses (score < 3) and raises the probability of very high-quality ones (score > 9), whereas offline DPO hardly changes the distribution at these extremes. Adding visual supervision signals shifts the overall distribution in a positive direction, though it makes little extra difference to the probabilities at the lowest and highest scores. These observations suggest that the benefits of online training and those of visual supervision are orthogonal, thereby explaining OViP’s stronger improvements achieved.
@@ 5.1 LVLM Hallucination
Works of synthetic data construction for mitigating hallucination in LVLMs can be broadly categorized into image-related synthesis and text-only synthesis. On the image side, several approaches leverage entity extraction and masking to perform targeted image editing, generating visually similar but semantically distinct counterfactuals (Xie et al., 2024; Lu et al., 2025). In contrast, HallusionBench (Guan et al., 2024) adopts a manual approach, carefully crafting counterfactual images to probe specific failure modes. Other works take a generative perspective: SynthVLM (Liu et al., 2024c) and SynthEmbedding (Sharifzadeh et al., 2024) utilize off-the-shelf models to synthesize new images or directly generate image embeddings for hallucination-aware training. Meanwhile, text-side data augmentation can also be used in LVLM training. VoCoT (Li et al., 2024) introduces new prompting patterns and response types to generate hallucination-prone QA data at scale. Other works such as Zhou et al. (2024a), Sarkar et al. (2024), Amirloo et al. (2024) introduce noise through perturbation, masking, or controlled corruption to simulate erroneous responses. More recent approaches (Xiao et al., 2025; Yu et al., 2024a) aim to detect and correct hallucinated content at varying levels of granularity, from token-level edits to full-sequence rewrites.
@@
5.2 Allocating More Computation on Training Sample Construction
Recent research has increasingly adopted the paradigm of allocating additional computation during training to get better training samples. Several studies utilize reinforcement learning with human or AI-generated feedback to guide VLM outputs. RLHF-V (Yu et al., 2024a) leverages fine-grained human annotations to correct hallucinated content, while RLAIF-V (Yu et al., 2024b) replaces human labels with feedback from ensembles of open-source models, significantly reducing annotation overhead. Similarly, OPA-DPO (Yang et al., 2025) employs an on-policy editing step prior to DPO, aligning training samples closely with model predictions to enhance data efficiency. CLIP-based methods dynamically filter self-generated samples for high-quality training pairs (Ouali et al., 2024; Zhou et al., 2024b). Other methods integrate auxiliary reward models or evaluators during training, providing continuous and adaptive feedback loops (Sun et al., 2024; Yan et al., 2024). Additionally, recent approaches incorporate reasoning or editing mechanisms directly into training, using iterative self-feedback or generative data augmentation techniques to dynamically refine model outputs (Zhao et al., 2023; Kim et al., 2024). These strategies improve model alignment and factuality by enriching the quality and relevance of supervision signals during training.
@@ 6 Conclusion
In this work, we propose the Online Vision-language Preference Learning (OViP) framework to efficiently address the hallucination problem in LVLMs. By integrating online preference learning with image-aware training, OViP enables real-time construction of high-quality contrastive data during training. Furthermore, to better assess the trade-offs between hallucination reduction and overall performance, we refine and extend existing evaluation protocols. Experimental results demonstrate that OViP significantly outperforms prior offline/online training approaches, achieving substantial hallucination reduction while preserving general vision-language capabilities, which many existing offline methods fail to preserve. Our investigation into training dynamics also sheds light on the underlying mechanisms behind OViP’s effectiveness.
Ethics Statement
Reproducibility Statement
We provide detailed descriptions of the training and evaluation setups in Appendix A and Appendix B. In addition, we include anonymized training and evaluation code, instructions for running the experiments, and information on accessing the relevant datasets in the supplementary materials.
@@ A.1.1 Metric Design
Let $a_{i},i\in\left\{1,2,3,4,5\right\}$ denotes $\mathrm{F1_{\mathrm{AMB-gen}},Score_{MMHal},F1_{\mathrm{ObjectHal}},LV_{\mathrm{score}},F1_{AMB-dis}}$ respectively, namely the results on each benchmark, superscript “base” represents performances of the baseline model and “ref” represents the set reference performances. Then HRI is calculated as:
@@ A.1.5 Further Study
A.1.6 Fairness
When aggregating different metrics through weighted averaging, it is necessary to account for the relative importance of each metric. Here, we define the potential improvement of a metric by considering its maximum observed gain in comparable experiments, and assign its weight as the inverse of this gain to normalize across metrics. For example, if metric A shows a maximum improvement of 2 points while metric B improves by 4 points, we assume that an equally strong model would, on average, achieve only half as much gain on A as on B. Consequently, each point of improvement on A should be considered twice as important as a point on B. Compared with simple averaging, this weighting scheme better reflects the relative significance of different metrics and is less prone to being gamed.
A.2 Benchmarks
  • MMHal-Bench (MMHal) (Sun et al., 2024) is a model-evaluated question-answering benchmark covering 8 categories and 12 topics. While the original evaluation strategy uses GPT-4 to judge model responses, a text-only model will introduce considerable judging-time hallucinations and errors, so gpt-4o-2024-05-13 is better for evaluation. (Amirloo et al., 2024).
  • AMBER generative (AMBgen(Wang et al., 2023) is a judging-model-free benchmark for the image description task, comprising 1,004 samples. Chair measures the object-level hallucination rate as the average precision of objects mentioned in the model’s descriptions, while Cover indicates the recall of objects. We observe a noticeable trade-off between these two metrics across various methods, where improvements in one often come at the expense of the other. To provide a more balanced and overall assessment, we introduce a new F1 score calculated as the harmonic mean of Chair and Cover.
  • Object HalBench (ObjectHal) (Rohrbach et al., 2018) evaluates object-level completeness and hallucination rates. The generation prompts are augmented from Yu et al. (2024a). Chairr denotes the response-level hallucination rate. We also introduce an object-level F1 metric to comprehensively measures the balance between hallucination and object coverage. Objects extraction is performed using gpt-4o-2024-05-13.
  • Llava-Bench-in-the-Wild (LV) (Liu et al., 2023) evaluates models’ visual abilities, using 60 open-ended questions grounded in 24 diverse images from real-world and abstract scenarios. The evaluation is conducted using gpt-4o-2024-05-13.
  • AMBER discriminative (AMBERdis(Wang et al., 2023) includes 14,216 ‘Yes/No” questions regarding objects in image. We use the F1 score as its metrics.
A.3 Bad Cases
A.3.1 MMHal
A.3.2 AMBER-generative & ObjectHal
@@
Figure 9: OPA-DPO fails to mention the man, a deficiency that is captured by Cover score but often overlooked in previous evaluations. “vehicles” is incorrectly identified as a hallucination word.
@@ B.2 Settings
Training
Both the 7B and 13B models are trained for a single epoch using a cosine learning rate schedule with a global batch size of 16. We set $\beta=\beta_{1}=\beta_{2}=0.1$ in Eq. 4 and Eq. 5. Learning rates are 1e-6 for 7B model and 5e-7 for 13B model.
@@
Our hyperparameter settings are based on preliminary experiments and empirical intuition. We observe that when the model assigns a score between 0 and 3, the responses tend to contain significant errors, while scores of 7 and above generally indicate correct answers. The more strict the preference filtering criteria is, the higher the data quality tends to be; however, this also leads to fewer preference pairs satisfying the condition. Therefore, our choice of hyperparameters is based on a balance among empirical observations, data quantity, and data quality.
@@ B.2.2 Ablation on Online Learning
B.3 Results
The training results for online and offline are shown in Table 5. Online training is significantly more effective in mitigating hallucinations.
Table 5: Online v.s. Offline detailed results
AMBgen MMHal ObjectHal LV AMBdis
[…]
Baseline 7.1 50.0 65.01 1.90 51.38 72.40 57.20 85.5
DPOonline 4.7 50.0 65.59 2.38 31.58 71.70 56.10 86.7
DPOoffline 7.0 48.3 63.58 2.06 52.61 72.55 53.60 85.9
OViPonline 4.0 51.1 66.70 2.52 33.22 73.50 63.10 87.1
OViPoffline 5.2 49.9 65.38 2.35 46.34 72.39 60.20 86.6
B.4 Training Dynamics
Figure 10: Sampling statistics during training. The blue curve shows the probability that, when sampling 16 responses with Temperature 1.2 for the same prompt, multiple identical responses appear (i.e., the number of distinct responses is fewer than 14). The red curve shows the average number of distinct responses obtained when sampling 16 times with Temperature 1.2.
Figure 11: Perplexity changes for IND and OOD sequences. The perplexity of base model’s generation is relatively low.
The model’s initially skewed output distribution leads to a high rate of duplicate samples (in Figure 11, Duplicate Response Rate surpasses 11.0% at first) and very low perplexity Figure 11 in the generated responses, which is not conducive to optimization. In the early stages of training, the output distribution gradually flattens, but the limited exploratory scope prevents the model from identifying the correct optimization direction, resulting in stagnation of performance metrics. As the coverage of the distribution expands, the model can effectively explore the correct update directions, allowing training to get on track and performance to accelerate.
@@ Appendix E Limitations
Appendix F Prompts for Judgment and Negative Image Generation
@@
Prompt for image generation-Part 1 # Task Given an unknown image-related question, a correct answer, and an inaccurate response, carefully analyze the differences between the response and the answer. Then, provide a brief description of the image so that it aligns with the correct answer and differs from the incorrect responses. In other words, infer the content of the image. # Example [Example 1] **<Question>** What is on the sandwich? **<Answer>** The sandwich has tomatoes and lettuce on it. **<Response>** The sandwich has a slice of egg and tomato on it. **<Output Description>** A sandwich with only tomatoes and lettuce on it. **Explanation**: The answer mentions lettuce and tomato, while the incorrect response mentions tomato and egg. So there is no egg on the sandwich. [Example 2] **<Question>** Can you point out the details that make this image unique? **<Answer>** In the image, there is a plate with a slice of pizza topped with tomatoes, herbs, and cheese. The distinctive detail about the image is that the pizza is missing two slices, leaving just one slice remaining on the plate. This suggests that someone has already started enjoying the pizza. The slice appears to be well-cooked and freshly served, creating an appetizing and mouthwatering scene for the viewer. **<Response>** In the image, there is a slice of pizza on a plate with tomatoes and cheese. The pizza appears to be homemade and has been cut into two pieces. The tomatoes are sliced in half, revealing their juicy interior. The cheese on top of the pizza is melted, creating a delicious-looking dish. Additionally, there is a fork nearby, suggesting that someone might be planning to enjoy this pizza soon. **<Output Description>** A plate with a one-third remaining piece of pizza, topped with herbs, cheese, and tomatoes; someone has finished eating and left. **Explanation**: The answer mentions that only one-third of the pizza remains and that someone has just finished eating and left, which is inconsistent with the response. Therefore, the image should include these two features.
Table 7: The prompt for image generation.
Prompt for image generation-Part 2 [Example 3] **<Question>** Bird or cow? **<Answer>** Bird **<Response>** The bird in the image is a small, brown and white bird with a distinctive head shape and coloration. It is not a cow. The bird is perched on a branch, which is situated in front of a white building. **<Output Description>** A big, blue bird perched on a branch in front of a black building. **Explanation**: Both the answer and the response mention the bird, but the response is more detailed. So the description should be contrastive to the features of the bird in the response. # Requirements - The description should be brief but precise. - If both the answer and the response are long, focus on describing the one or two most significant differences. - Do not provide any analysis or explanation; only describe the image. - A common approach is to describe what is present in the image and what is missing. **<Question>** question **<Answer>** answer **<Response>** response **<Output Description>**
Table 8: The prompt for image generation.